Suppose you invested $1,000 in stocks 5 years ago and your account is now worth $2,150. Please calculate the annual rate of return on this investment
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

2150=1000*(1+r/100)^5

(2150/1000)^(1/5)=(1+r/100)

(1+r/100)=1.1654

r=1.1654-1

=16.54%(Approx)